What is the center and the radius of the circle represented by the equation :
x^2+(y-4)^2=25

Answer by scott8148(6628)

the general equation for a circle is (x-h)^2+(y-k)^2=r^2 ... this is a circle centered at (h,k) with a radius of r

in your case, (x-0)^2+(y-4)^2=5^2 ... so center is (0,4) and radius is 5